# Towards Better Open-Ended Text Generation: A Multicriteria Evaluation Framework

### v2

@@ 7 Conclusion
In this paper, we addressed the pressing need for more reliable evaluation methodologies in open-ended text generation by proposing a novel benchmarking framework grounded in multicriteria analysis. Traditional metrics, such as coherence, diversity, generation perplexity, and MAUVE (see Garces-Arias et al. (2024b); Su and Collier (2023)), while considered separately, often yield inconsistent to suboptimal results. Furthermore, they have demonstrated a limited correlation with human judgment. We introduced Q*Text, a new automatic evaluation metric designed to offer a more robust measure of generated text quality.
To ensure robust and statistically valid conclusions, we adopted recent advancements in multicriteria benchmarking theory, specifically leveraging partial order depth functions to avoid the limitations of conventional ranking methods. Instead of relying on either Pareto-fronts or weighted aggregations, which are subject to biases or arbitrary weighting, our approach embraces incomparability. This allows us to generate partial orders of competing decoding strategies, offering a more nuanced and meaningful analysis of performance across multiple criteria.
By applying these methodologies in extensive experiments across diverse datasets and models, we have shown that our proposed approach not only enhances the interpretability of text generation evaluations but also provides a practical, data-driven framework for benchmarking algorithms in this domain. We hope that these insights will contribute to a more rigorous evaluation of text generation models and inspire further research in the field of multicriteria benchmarking strategies. Summing up, this work makes the following contributions.
Contributions
  1. 1.
    We employ novel benchmarking techniques for multicriteria evaluation to identify optimal combinations of model architectures and decoding strategies.
  2. 2.
    In order to identify which benchmarking technique is best for which practical use case, we discover two main benchmarking scenarios – a distinction that might be of indepedent interest for future research.
  3. 3.
    We introduce a new summarizing metric, Q*Text, which consolidates the strengths of existing automatic evaluation metrics, facilitating the ranking of decoding methods.
  4. 4.
    We extend the evaluation dataset by Garces-Arias et al. (2024a), by including sentence-level metric computation and incorporating generation perplexity to facilitate ranking of decoding methods.
  5. 5.
    We critically discuss benchmarking methods for evaluating models and decoding strategies with clear recommendations for practitioners (Section 8).
  6. 6.
    We make our efficiently (see Appendix A.7) implemented benchmarking methods and datasets publicly accessible at this link.
@@ 8 Key Takeaways and Practical Recommendations
    Is it sufficient to rank methods, or is metric information about the methods’ performances required? (Scenario 1 and 2 in Section 1)
@@
    Does the use case require a total or partial ordering method, i.e., should the evaluation allow for incomparability among some methods, or should it enforce comparability of all methods? (Section 4)
In case metric information is required and comparability of all methods should be enforced, we recommend our novel aggregation metric Q*Text, see Section 5. If the metric information is not the overall aim, but comparability should still be enforced, we recommend using the Bradley-Terry model, see Section 4.1. Eventually, if a ranking is required that allows for incomparability, we recommend deploying ufg-depth; see Section 4.2.
@@
While our study presents three different benchmarking approaches, this by no means covers the full range of different benchmarking strategies that aim to address the different objectives, i.e. selecting an estimated best method vs. estimating the performance structure of methods. Therefore, this article provides only a glimpse of the complexity and different approaches to multi-metric evaluation.
Besides this, further limitations merit attention. First, our experiments focused on a limited set of decoding strategies and language models. Alternative methods—such as contrastive decoding (Li et al., 2023), typical sampling (Meister et al., 2023), and adaptive contrastive search (Garces-Arias et al., 2024b)—were not analyzed and may provide insights that refine or challenge our findings.
Secondly, the choice of metrics is debatable. Our reliance on model-dependent metrics—such as coherence, which is measured by an ideally unbiased OPT 2.7B model Zhang et al. (2022), raises questions about their robustness across different models and datasets. Moreover, including further metrics might enhance the robustness and generalizability of our conclusions.
Thirdly, also including more diverse datasets could enhance the robustness and generalizability of our conclusions.
Additionally, while our work focuses on open-ended text generation, the methodologies and insights may apply to other NLP tasks like summarization and machine translation, which present different challenges and evaluation criteria. Applying our framework to these tasks could provide valuable perspectives on evaluation metrics and benchmarking strategies in broader contexts.
@@ Acknowledgments
Currently not added due to anonymization purposes.
Appendix A Appendix
A.1 Automatic metrics
@@
A.2 Q*Text
Q*Texts results from the harmonic mean of normalized and penalized coherence, diversity, and generation perplexity (see Section A.1).
A.2.1 Normalization
Normalized Coherence:
Originally ranging from $\infty$- ∞ to 0, with higher values indicating better coherence. We applied min-max normalization based on the observed minimum ($C_{\min}$) and maximum ($C_{\max}$) coherence values in our dataset:
$C_{\text{norm}}=100\times\frac{C-C_{\min}}{C_{\max}-C_{\min}}$
Normalized Diversity:
Naturally ranging from 0 to 100, with higher values indicating greater diversity. Since it already fits our desired scale, no transformation was necessary:
$D_{\text{norm}}=D$
Normalized Perplexity:
Ranging from 1 to $\infty$, with lower values indicating better performance. We first inverted the perplexity to align the orientation (higher is better) and then applied min-max normalization:
$P^{\prime}=\frac{1}{P}$
$P_{\text{norm}}=100\times\frac{P^{\prime}-P^{\prime}_{\min}}{P^{\prime}_{\max}-P^{\prime}_{\min}}$
A.2.2 Penalization
While higher values in each normalized metric generally indicate better performance, excessively high values can lead to undesirable effects in generated text, such as repetitiveness or incoherence. To mitigate this, we applied a sigmoid-based penalization function to each normalized metric, discouraging values that are too close to the maximum:
Penalized $\text{Penalized Score}=S_{\text{norm}}\times\left(1-\frac{1}{1+e^{-k(S_{\text{norm}}-t)}}\right)$
where $S_{\text{norm}}$ is the normalized score (either $C_{\text{norm}}$, $D_{\text{norm}}$, or $P_{\text{norm}}$), $t$ is the threshold beyond which penalization starts, $k$ is the steepness of the penalization curve. The function $\frac{1}{1+e^{-k(S_{\text{norm}}-t)}}$ yields a value between 0 and 1, increasing sharply as $S_{\text{norm}}$ exceeds $t$. We empirically found a balanced penalization with $t=95$ and $k=5$.
This penalization ensures that as a metric approaches excessively high values, its contribution to Q*Text diminishes smoothly, preventing any single metric from disproportionately influencing the aggregated score.
A.2.3 Aggregation
After normalization and penalization, we combined the adjusted scores into a single Q*Text value using the harmonic mean. The harmonic mean is particularly suitable when averaging ratios or rates and when we want to ensure that low values in any component significantly impact the overall score:
[…]
where: $n$ is the number of metrics being combined (in this case, 3), $S_{\text{penalized},i}$ represents each of the penalized scores (Penalized Coherence, Penalized Diversity, Penalized Perplexity).
A.3 Result of pairwise comparison
@@
The following tables consider the pairwise comparisons of the methods on the generation level, e.g., we count on how many generations one method strictly outperforms another method, compared to Section 4.1. Since we are comparing $354$ many methods (consisting of model and decoding strategy combination), we have to consider $354\cdot 353=124962$ many pairwise comparisons.
@@
Table 4: All pairwise comparisons of two methods where Method 1 strictly dominates Method 2 based on the three metric perplexity, coherence and diversity on all 1314 generations of WikiText-103. Count denotes the number of generations where Method 1 strictly dominates Method 2.
Table 5 collects all pairwise comparisons where Method 1 strictly dominates Method 2 based on all 2000 generations of Wikinews and the metrics perplexity, diversity and coherence. Moreover, we can observe that for 878 of all 124962 pairwise comparisons we have that at least on 90% of the generations method 1 dominates method 2 strictly. For 25108 pairwise method comparisons, we obtain that method 1 never strictly dominates method 2 (i.e. on every generation method 2 either dominates method 1 or the three metrics disagree on the dominance structure or a completely equal).
@@
When we merge the three datasets WikiText-103, Wikinews and Book, we consider $1314+2000+1947=5261$ generations and 124962 pairwise comparisons based on each generation. Comparing the tables 4, 5, 6 we find that there is no pairwise comparison that occurs in each table. Therefore, there is no pair of two methods where method 1 dominates method 2 based on all 5261 generations. With 4601 is the dominance of Mistral 3 CS ((’0.8’, ’10’)) over GPT2-XL CS ((’1.0’, ’10’)) the one that occurs most often. For 2990 pairwise comparison at least on 90% of the generations method 1 dominates method 2 strictly. In 9191 pairwise method comparisons, we obtain that method 1 never strictly dominates method 2 (i.e. on every generation method 2 either dominates method 1 or the three metrics disagree on the dominance structure or a completely equal).
A.4 Results of the extended Bradley-Terry model
In this section, we present the complete result of the extended Bradley-Terry model for all $354$ methods. Note that the higher the estimated worth parameter of the extended Bradley-Terry model, the higher the estimated probability that the method outperforms another method. Hence, the method with the highest worth parameter is, according to the extended Bradley-Terry model, the one that outperforms all others.
For reasons of clarity and comprehensibility, we decided to show here only a snippet, but the full result can be easily and fast obtained by the already stored results in GitHub-repository. Table 7 denotes the worth parameter based on WikiText-103, Table 8 on Wikinews, Table 9 on Books and all three datasets combined can be seen in Tables 10 and 11. All computations are based on the metrics of perplexity, coherence, and diversity.
@@ A.4 Results of the extended Bradley-Terry model
Table 9: Estimated worth parameter of the extended Bradley-Terry model based on Book dataset and the metric coherence, diversity and perplexity.
Method Estimated worth parameter
@@
Method Estimated worth parameter
@@
Table 10: Estimated worth parameter of the extended Bradley-Terry model based on WikiText-103, Wikinews, and Book datasets together and the metric coherence, diversity, and perplexity (1/2).
Method Estimated worth parameter
@@
Method Estimated worth parameter
@@
A.5 Discussion of the Ufg-depth Results
@@
A.6 Results of Q*Text
Based on the Q*Text metric introduced in Section 5, we can induce a total ordering of decoding methods. Tables 12, 14, 16 and 18 illustrate the results for the most dominant decoding models, strategies, hyperparameters and methods, respectively. On the other hand, We observe in Tables 13, 15, 17 and 19 the results for the least dominant decoding models, strategies, hyperparameters and methods.
@@ A.6 Results of Q*Text
Figure 5: Decoding methods with the largest rank discrepancies between the extended Bradley-Terry model and Q*Text. Here, the extended Bradley-Terry model notably favors low-diversity methods, such as beam search, while Q*Text tends to rank highly diverse methods higher. This highlights the differing emphases of each approach on diversity in decoding strategies.
@@
Examples of best and worst decoding methods according to Q*Text
In this part, we present a case study comparing various decoding strategies based on their performance as evaluated by Q*Text.
Prompt "Headlam served as Officer Commanding North @-@ Western Area in 1946, and as Director of Training from 1947 to 1950. In 1950 - 51,", Q*Text
Deepseek CS (0.6, 15) "he was Chief of Staff (C.O.S.) of the 1st Parachute Brigade, and in 1951 2̆013 52, he was G.O.C. of the 1st Parachute Brigade. He was promoted to Brigadier… (text continues)" 80.16
GPT2-XL CS (1.0, 20) "Headless Horsemurdertrud (his nom de v0̆0eal et re, a translators and transcriph.coastavarism in Latin-Ceux du Ponto-Monahawak at Keremej, New Heythorhampsthembryghtes-sur Meriont Bay of Sulaunkhuyut/Arap Moheluwempeh is one hundred nautilus middelveth the seagoch that swim to an anchord-like… (text continues)" 0.02
Qwen 2 CS (1.0, 20) " Headelm oversa 4̆e86 Operation Agulhene535af69, to reoccupyo portug 1 st colonial outst a ncesin AFRicacoun trie sto preva l sepa rabli st mov me s . In th es on he wro 4he couthoea2 rsh ipp ublicatin hitherto re t aine o secret… (text continues)" 0.02
Llama 3 beam (3) "he served as Commanding Officer of the 1st Battalion, The Queen’s Royal Regiment (West Yorkshire Regiment). In 1951 2013 52, he served as Commanding Officer of the 2nd Battalion, The Queen’s Royal Regiment (West Yorkshire Regiment). In 1952 2013 53, he served as Commanding Officer of the 3rd Battalion, The Queen’s Royal Regiment (West Yorkshire Regiment). In 1953 2013 54, he served as Commanding Officer of the 4th Battalion, The Queen’s Royal Regiment … (text continues)" 0.02
Table 20: Case Study: Comparison of multiple decoding methods for a prompt from the Wikitext corpus. The first five rows show examples generated by high-ranked methods, while the next three rows display those from low-ranked methods. Human-generated reference text is included for comparison.
A.7 Computation Complexity of the Algorithms
Here, we give some information about the complexity of the above algorithms. Therefore, let $n$ be the number of methods we compare to $m$ generations.
The extended Bradley-Terry model and the union-free generic depth both require that we go through all pairwise comparisons for all generations. This has a complexity of $\mathcal{O}(n^{2}m)$. For the extended Bradley-Terry model, we need to compute a generalized linear model. The definition of the corresponding design matrix also requires $\mathcal{O}(n^{2}m)$. Computing the union-free generic depth can be very time-consuming. In the worst case, we have to go through every subset of the partial orders, which gives a worst-case scenario of $\mathcal{O}(2^{m})$. However, in our situation (and many many others) it can be shown that it is sufficient to consider only all subsets of cardinality 10 or lower. Moreover, by using the connectedness property, see Blocher et al. (2024), we can further reduce the computation time. For a detailed discussion of the computational time of union-free generic depth see Blocher et al. (2024).

### v3

@@ 7 Conclusion
In this work, we analyze the challenge of evaluating open-ended text generation by introducing a multicriteria benchmarking framework that supports both relative and absolute rankings of decoding methods. We present three complementary approaches—the extended Bradley-Terry model, the union-free generic (ufg) depth, and Q*Text, a unified metric that harmonizes coherence, diversity, and perplexity into a single score. Moreover, we show that our framework captures nuanced trade-offs among metrics and avoids misleading comparisons when methods excel on different criteria.
Extensive experiments involving six large language models, three distinct domains (news, Wikipedia, stories), and over 1.8 million generated continuations demonstrate the practical benefits of our approach. The extended Bradley-Terry model yields interpretable “worth” parameters that reflect overall preference probabilities, while ufg-depth uncovers central and atypical ranking structures, highlighting when decoding methods are genuinely incomparable. Q*Text further enables direct comparison and quantification of performance gaps, revealing that balanced hyperparameter settings outperform extreme configurations and that smaller models can rival larger ones under appropriate decoding choices. Taken together, these contributions provide practitioners and researchers with a more reliable, data-driven basis for selecting and designing decoding methods in open-ended text generation, paving the way for more holistic benchmarking practices.
@@ 8 Key Takeaways and Practical Recommendations
    Is it sufficient to rank methods, or is metric information about the methods’ performances required? (Scenario 1 and 2 in §1)
@@
    Does the use case require a total or partial ordering method, i.e., should the evaluation allow for incomparability among some methods, or should it enforce comparability of all methods? (§4)
In case metric information is required and comparability of all methods should be enforced, we recommend our novel aggregation metric Q*Text, see §5. If the metric information is not the overall aim, but comparability should still be enforced, we recommend using the Bradley-Terry model, see §4.1. Eventually, if a ranking is required that allows for incomparability, we recommend deploying ufg-depth; see §4.3.
@@
While our study presents three different benchmarking approaches, this by no means covers the full range of different benchmarking strategies that aim to address the different objectives, i.e., selecting an estimated best method vs. estimating the performance structure of methods. Therefore, this article provides only a glimpse of the complexity and different approaches to multi-metric evaluation.
Besides this, further limitations merit attention. First, our experiments focused on a limited set of decoding strategies and language models. Alternative methods—such as contrastive decoding (Li et al., 2023), typical sampling (Meister et al., 2023), and adaptive contrastive search (Garces Arias et al., 2024)—were not analyzed and may provide insights that refine or challenge our findings.
Secondly, the choice of metrics is a matter of debate. Our reliance on model-dependent metrics, such as coherence, which is measured by an ideally unbiased OPT 2.7B model Zhang et al. (2022), raises questions about their robustness across different models and datasets He et al. (2023). Moreover, including further metrics might enhance the robustness and generalizability of our conclusions.
Additionally, while our work focuses on open-ended text generation, the methodologies and insights may also apply to other NLP tasks, such as summarization and machine translation, which present different challenges and evaluation criteria. Applying our framework to these tasks can provide valuable insights into evaluation metrics and benchmarking strategies in broader contexts.
@@ Acknowledgments
Hannah Blocher received financial support via a stipend from Evangelisches Studienwerk Villigst e.V. Julian Rodemann acknowledges support by the Federal Statistical Office of Germany within the co-operation project "Machine Learning in Official Statistics" as well as by the Bavarian Institute for Digital Transformation (bidt) and the Bavarian Academy of Sciences (BAS) within a graduate scholarship. Matthias Aßenmacher received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of BERD@NFDI, under grant number 460037581.
Appendix
Appendix A Automatic metrics
@@
Appendix B Union-Free Generic Depth
General definitions.
Let $M$ be a set of items/models. $p\subseteq M\times M$ is a partial order (poset) iff $p$ is reflexive (i.e. for all $m\in M,(m,m)\in p$ ), transitive (i.e. $\left.\left(m_{1},m_{2}\right),\left(m_{2},m_{3}\right)\in p\Rightarrow\left(m_{1},m_{3}\right)\in p\right)$ and antisymmetric (i.e. $\left(m_{1},m_{2}\right),\left(m_{2},m_{1}\right)\in p\Rightarrow m_{1}=m_{2}$ ). A closure operator on a set $\Omega$ is a function $\gamma:2^{\Omega}\rightarrow 2^{\Omega}$ that is extensive (i.e. for all $A\subseteq\Omega$ we have $A\subseteq\gamma(A)$ ), increasing ( $A\subseteq B\subseteq\Omega\Rightarrow\gamma(A)\subseteq\gamma(B)$ ) and idempotent (for all $A\subseteq\Omega,\gamma(A)=\gamma(\gamma(A)))$
Union-free generic depth.
The definition of the ufg-depth, see Blocher et al. (2024), is analogous to the definition of the simplicial depth on $\mathbb{R}^{d}$, see Liu (1990). Hence, we first have to consider a closure operator $\gamma:2^{\mathcal{P}}\rightarrow 2^{\mathcal{P}},P\mapsto\left\{p\in\mathcal{P}\mid\cap_{\tilde{p}\in P}\>\tilde{p}\subseteq p\subseteq\cup_{\tilde{p}\in P}\>\tilde{p}\right\}$. Then a poset $p\in\mathcal{P}$. This is indeed a closure operator and now can be used to generalize the notion of $d+1$ simplices. As described above, we therefore define the set
$\mathcal{S}=\{P\subseteq\mathcal{P}\mid\text{ Condition }(C1)\text{ and }(C2)\text{ hold }\}$
with conditions (C1) $P\subsetneq\gamma(P)$ and (C2) there does not exist a family $(\tilde{P}i)i\in 1,\ldots,\ell$ such that for all $i\in 1,\ldots,\ell\tilde{P}i\subsetneq P$ and $\bigcup_{i\in 1,\ldots,\ell}\gamma(\tilde{P}i)=\gamma(P)$. Note, the (empirical) ufg-depth is given by: Let $p_{1},\ldots,p_{n}\in\mathcal{P}$ be a sample with corresponding empirical probability measure $\nu_{n}$ (equipped with the power set as $\sigma$-field). Then, the (empirical) union-free generic ( $ufg$) depth is given by
$D_{n}(p)=\begin{cases}0,&\text{if }\forall S\in\mathcal{S}:\prod_{\tilde{p}\in S}\nu_{n}(\tilde{p})=0\\[3.00003pt] c_{n}\displaystyle\sum_{\begin{subarray}{c}S\in\mathcal{S}\\ p\in\gamma(S)\end{subarray}}\prod_{\tilde{p}\in S}\nu_{n}(\tilde{p}),&\text{else}\end{cases}$
with $c_{n}=\left(\sum_{S\in\mathcal{S}}\prod_{\tilde{p}\in S}\nu_{n}(\tilde{p})\right)^{-1}$. Note that since $\nu_{n}(p)=0$ if $p\in\mathcal{P}$ is not observed, we can restrict the set $\mathcal{S}$ to $\mathcal{S}_{\text{obs }}=\left\{S\in\mathcal{S}\mid S\subseteq\left\{p_{1},\ldots,p_{n}\right\}\right\}$ consisting only of the observed posets.
Example: As example consider the four methods Mistral 3 $\mathrm{CS}((0.6,15))$ (here denoted as $\left.m_{1}\right)$, Mistral 3 $\mathrm{CS}((0.4,3))$ (here denoted as $\left.m_{2}\right)$, Mistral 3 $\mathrm{CS}((0.8,3))$ (here denoted as $m_{3}$ ), and Mistral 3 $\mathrm{CS}((0.4,20))$ (here denoted as $m_{4}$ ). Assume that the quality metrics provide us with the following four posets:
Let $S=\{(m_{i},m_{i})\mid i\in\{1,2,3,4\}\}$. Then:
$p_{1}$ $=S\cup\{(m_{1},m_{2})\}$
$p_{2}$ $=S\cup\{(m_{1},m_{3})\}$
$p_{3}$ $=S\cup\{(m_{1},m_{2}),(m_{2},m_{3}),(m_{1},m_{3})\}$
$p_{4}$ $=S\cup\{(m_{1},m_{4})\}$
Then, with the closure operator above, we get that $p_{3}\notin\gamma\left(p_{1},p_{2}\right)$ (note that also incomparabilities are of interest via the union in the definition of the closure operator). The set $\mathcal{S}_{\text{obs}}=\{\{p_{1},p_{2}\},\{p_{1},p_{4}\},\{p_{2},p_{4}\},\{p_{3},p_{4}\},$$\{p_{1},p_{2},p_{3}\},\{p_{1},p_{2},p_{4}\},\{p_{2},p_{3},p_{4}\}\}$. With this, the ufg-depth of $D_{n}\left(p_{1}\right)=6/7$ and $D_{n}\left(p_{4}\right)=5/7$. Hence, $p_{1}$ is more central than $p_{4}$.
Appendix C Results of Pairwise Comparisons
@@
The following tables consider the pairwise comparisons of the methods on the generation level, e.g., we count on how many generations one method strictly outperforms another method, compared to §4.1. Since we are comparing $354$ many methods (consisting of model and decoding strategy combination), we have to consider $354\cdot 353=124962$ many pairwise comparisons.
@@
Table 5: All pairwise comparisons of two methods where Method 1 strictly dominates Method 2 based on the three metric perplexity, coherence, and diversity on all 1314 generations of WikiText-103. Count denotes the number of generations where Method 1 strictly dominates Method 2.
Table 6 collects all pairwise comparisons where Method 1 strictly dominates Method 2 based on all 2000 generations of Wikinews and the metrics perplexity, diversity, and coherence. Moreover, we can observe that for 878 of all 124,962 pairwise comparisons we have that at least on 90% of the generations method 1 dominates method 2 strictly. For 25,108 pairwise method comparisons, we obtain that method 1 never strictly dominates method 2 (i.e., on every generation, method 2 either dominates method 1 or the three metrics disagree on the dominance structure or are completely equal).
@@
When we merge the three datasets WikiText-103, Wikinews and Book, we consider $1314+2000+1947=5261$ generations and 124962 pairwise comparisons based on each generation. Comparing the tables 5, 6, 7 we find that there is no pairwise comparison that occurs in each table. Therefore, there is no pair of two methods where method 1 dominates method 2 based on all 5261 generations. With 4601 is the dominance of Mistral 3 CS ((’0.8’, ’10’)) over GPT2-XL CS ((’1.0’, ’10’)) the one that occurs most often. For 2990 pairwise comparison at least on 90% of the generations method 1 dominates method 2 strictly. In 9191 pairwise method comparisons, we obtain that method 1 never strictly dominates method 2 (i.e. on every generation method 2 either dominates method 1 or the three metrics disagree on the dominance structure or a completely equal).
Appendix D Results of the extended Bradley-Terry model
In this section, we present the complete result of the extended Bradley-Terry model for all $354$ methods.
@@ Appendix D Results of the extended Bradley-Terry model
Note that the higher the estimated worth parameter of the extended Bradley-Terry model, the higher the estimated probability that the method outperforms another method. Hence, the method with the highest worth parameter is, according to the extended Bradley-Terry model, the one that outperforms all others.
@@
For reasons of clarity and comprehensibility, we decided to show here only a snippet, but the full result can be easily and fast obtained by the already stored results in GitHub-repository. Table 8 denotes the worth parameter based on WikiText-103, Table 9 on Wikinews, Table 10 on Books and all three datasets combined can be seen in Table 11. All computations are based on the metrics of perplexity, coherence, and diversity.
@@
Table 10: Estimated worth parameter of the extended Bradley-Terry model based on Book dataset and the metric coherence, diversity, and perplexity.
Method Estimated
worth parameter
@@
Method Estimated
worth parameter
@@
Method Estimated
worth parameter
@@
Method Estimated
worth parameter
@@
Appendix E Discussion of the Ufg-depth Results
@@
Appendix F Results of Q*Text
Based on the Q*Text metric introduced in §5, we can induce a total ordering of decoding methods. Tables 12, 14, 16 and 18 illustrate the results for the most dominant decoding models, strategies, hyperparameters and methods, respectively. On the other hand, We observe in Tables 13, 15, 17 and 19 the results for the least dominant decoding models, strategies, hyperparameters and methods.
@@ Appendix F Results of Q*Text
Figure 4: Decoding methods with the largest rank discrepancies between the extended Bradley-Terry model and Q*Text. Here, the extended Bradley-Terry model notably favors low-diversity methods, such as BS, while Q*Text tends to rank highly diverse methods higher. This highlights the differing emphases of each approach on diversity in decoding strategies.
@@
Appendix G Q*Text Hyperparameters
Line Pseudocode: Q*Text Hyperparameter Tuning
Input: Perplexity, Coherence and Diversity scores (P, C, D)
1 P_norm = (max(P) - P) / (max(P) - min(P))
2 C_norm = (C - min(C)) / (max(C) - min(C))
3 D_norm = (D - min(D)) / (max(D) - min(D))
4 $\theta$ = [1,1,1,0.5,0.5,0.5,1,1,1]
5 bounds_w = [[0.1,5],[0.1,5],[0.1,5]]
6 bounds_$\mu$ = [[0,1],[0,1],[0,1]]
7 bounds_$\alpha$ = [[0.1,10],[0.1,10],[0.1,10]]
8 for trial in range(max_trials):
9    $\theta$_new = $\theta$ + random_normal(0, 0.1)
10    $\theta$_new = clip($\theta$_new, bounds)
11    for i in range(N):
12     penalty_p = exp(-$\alpha_{1}$(P_norm[i]-$\mu_{1}$)2)
13     penalty_c = exp(-$\alpha_{2}$(C_norm[i]-$\mu_{2}$)2)
14     penalty_d = exp(-$\alpha_{3}$(D_norm[i]-$\mu_{3}$)2)
15     QText[i] = ($w_{1}$_norm[i]penalty_p +
16         $w_{2}$_norm[i]penalty_c +
17         $w_{3}$_norm[i]penalty_d) / ($w_{1}$+$w_{2}$+$w_{3}$)
18    $\rho$ = spearman_corr(QText, Human)
19    if $\rho$ > best_$\rho$: $\theta$_best = $\theta$_new
20 return $\theta$_best
Table 20: Q*Text Optimization Algorithm
Algorithm explanation: Lines 1-3 normalize metrics to [0,1]. Lines 5-7 define parameter bounds for weights ($w_{i}\in[0.1,5.0]$), targets ($\mu_{i}\in[0.0,1.0]$), and penalties ($\alpha_{i}\in[0.1,10.0]$), this bound definition aims at (i) preventing zero weights while allowing one metric to dominate, (ii) match the normalized metric range, and (iii) ensure positive penalties with reasonable strength. Lines 9-10 perturb parameters with Gaussian noise and clip to bounds. The optimization maximizes Spearman correlation $\rho$ with human ratings.
Parameter Symbol Value
Metric Weights
Perplexity Weight $w_{1}$ 0.586
Coherence Weight $w_{2}$ 0.834
Diversity Weight $w_{3}$ 3.853
Gaussian Target Values ($\mu$)
Perplexity Target $\mu_{1}$ 0.458
Coherence Target $\mu_{2}$ 0.000
Diversity Target $\mu_{3}$ 0.854
Gaussian Penalty Strength ($\alpha$)
Perplexity Penalty $\alpha_{1}$ 2.579
Coherence Penalty $\alpha_{2}$ 1.496
Diversity Penalty $\alpha_{3}$ 7.370
Table 21: Optimal Q*Text Hyperparameters (Spearman $\rho_{s}=0.5545$)
Parameter Interpretation.
The optimized parameters reveal insights about text quality assessment.
Diversity dominance: The substantially higher weight for diversity ($w_{3}=3.853$) compared to perplexity ($w_{1}=0.586$) and coherence ($w_{2}=0.834$) indicates that lexical variety is the most discriminative factor for human preferences in our dataset.
Target preferences: The optimal targets suggest humans prefer moderate perplexity levels ($\mu_{1}=0.458$), minimal coherence constraints ($\mu_{2}=0.000$), and high diversity ($\mu_{3}=0.854$).
Penalty sensitivity: The high diversity penalty strength ($\alpha_{3}=7.370$) enforces strict adherence to the diversity target, while the moderate perplexity penalty ($\alpha_{1}=2.579$) and lenient coherence penalty ($\alpha_{2}=1.496$) allow more variation in these two dimensions.
Figure 5: Correlation between Q*Text scores and human ratings across six text generation methods. Each point represents a text sample, colored by generation method. The dashed line shows the linear regression fit. Q*Text achieves a moderate positive correlation (Spearman $\rho=0.5545$, $p<0.001$) with human evaluations, demonstrating its effectiveness in capturing human preferences for text quality.